# Deep Support Vectors

### v1

@@ 4.2 Deep Support Vectors, From Dust to Glory
Firstly, the primal feasibility condition which is equivalent to Eq. (4) questions the fundamental performance of the model. That is, the trained model $\theta^{*}$ must be able to correctly discriminate its support vectors. As presented in Sec. 4.1, instead of Eq. (4), we use a surrogate function for the loss:
$L_{i}=\begin{cases}0&\text{if }\operatorname*{argmax}_{c}\Phi_{c}(x_{i};\theta^{*})=y_{i},\\ L(\Phi(x_{i};\theta^{*}),y_{i})&\text{otherwise}.\end{cases}$
This is designed to match the primal condition by mimicking the Hinge loss. When each DSV is correctly classified, no loss is incurred. Otherwise, we adjust the DSVs to align them with the correct target, effectively optimizing the support vectors. This approach also implicitly enforces the complementary slackness condition, as $L_{i}$ does not promote increased confidence in the incorrect classification. Consequently, the final loss, which refines the support vectors, is computed as follows:
$L_{\text{primal}}=\frac{1}{n}\sum_{i=1}^{n}L_{i}.$ (7)
Here, $L$ can be any loss function and we have employed the cross-entropy loss in our experiments.
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L_{\text{stat}}=D(\theta^{*},-\sum_{i=1}^{n}\lambda_{i}\nabla_{\theta}L(\Phi(x_{i};\theta^{*}),y_{i}))}$ (8)
@@
Moreover, as shown in Algorithm 1, we satisfy the dual condition by ensuring the Lagrange multipliers $\lambda_{i}$’s are greater than zero and disqualify any $x_{i}$’s from being a support vector candidate if during optimization $\lambda_{i}$ becomes less than zero. The condition that is not directly satisfied is the complementary slackness. To directly fulfill the functional boundary for support vectors as specified in Sec. 3.2, we would need to be able to calculate the distance between functions, which is not only abstract but also requires a second-order computation cost. Therefore, we adopted a relaxed version of the KKT conditions that excludes this requirement. Furthermore, as demonstrated in Sec. 5.2, we have shown that DSVs implicitly satisfy the complementary slackness condition. This implies that DSVs meet every condition introduced in conventional SVM.
Figure 2: Conceptual overview of our algorithm. Initially sampled from noise, the DSVs evolve to intersect with the KKT solution subspace and the data manifold.
@@
In addition to satisfying the conventional KKT conditions, we have adopted additional criteria for determining DSVs. As much as determining the data distribution in a high-dimensional space, giving additional constraints will allow DSVs to approach the desired data manifold. The DeepKKT conditions form an DSV solution subspace as inferred in Fig. 2, however this does not guarantee that it is on the data manifold that the data is trained on. Among the DSVs that satisfy the DeepKKT conditions, we must select the ‘good’ DSVs closest to the manifold. Since the model is trained on the data manifold, it follows that the DSVs would also be formed within this manifold. By selecting DSVs that are in the intersection of the solution subspace and the data manifold, it will be possible to properly represent both the model and the training dataset.
To extract DSVs from the manifold, we leverage the principles of Self-Supervised Learning (SSL). SSL operates under the assumption that a robust model should exhibit augmentation invariance. Inverting this idea, we propose that, when evaluating images within the manifold, an effective model should assess them in an augmentation-invariant fashion. Put differently, the DSVs we identify should remain as DSVs even after augmentation.
@@ 5 Experiments
Figure 3: Illustration of entropy change with DSV candidates: We utilized DeepKKT conditions in the selection of DSVs. As training progresses, entropy increases, i.e., the DSVs move toward the decision boundary.
5.1 Implementation Details
We chose the ConvNet architecture (Goodfellow et al., 2016) for pretraining $\Phi(\cdot;\theta)$ on the SVHN dataset (Netzer et al., 2011), a digit dataset with dimensions similar to CIFAR-10 (Krizhevsky, 2009). To synthesize DSVs, we used a large learning rate of 0.1 and a $\rho$ value of 0.0001 were utilized for the SAM optimizer to adhere to the principle of margin maximization (Wang et al., 2023; Foret et al., 2020). For $L_{\text{stat}}$, we set $\alpha_{x}=10^{-3}$, $\alpha_{\lambda}=0.1$, and $\gamma=5$, respectively. For synthesis, it took 15 minutes in a single 1080TI GPU. We leave additional details in Appendix Sec.B
5.2 DSVs: Decoding Black-Box in Deep Learning
DSVs Implicitly Meet Complementary Slackness
@@ 5.2 DSVs: Decoding Black-Box in Deep Learning
As we discussed in Sec. 3.2, the principle of complementary slackness within the KKT conditions suggests that support vectors should be situated on the decision boundary. This aspect was not incorporated explicitly in our model due to computational costs and ambiguity. However, as indicated in Fig. 3, it appears that DSVs implicitly fulfill this condition. Notably, during the training process, we observe an increase in the entropy of DSV candidates.
From a probabilistic perspective, high entropy indicates greater uncertainty. This is a strong indicator that these samples are positioned near the decision boundary, as they are the ones for which the model is least certain about their classification. This aligns with the complementary slackness principle, which posits that support vectors play a crucial role in defining the decision boundary.
Predicting Model Dynamics With DSVs’ Inherent Importance
Figure 4: Correlation between the classwise mean test accuracy versus the sum of $\lambda$’s for training samples belonging to each class. Here, $\rho$ represents the Pearson correlation coefficient.
In SVM, the Lagrangian multiplier $\alpha$, acts as a natural measure to determine which support vector is more important than others. Similarly, DSVs can act a comparable role. Fig. 4 indicates a strong correlation between the sum of $\lambda$’s per each class and actual classwise test accuracy. This phenomenon is intriguing because, despite the model achieving nearly 100% accuracy during training due to overparameterization, DSVs can provide insights into the categorical generalization in the test phase.
Figure 5: Classification result of a mixed image of an original image and a DSV: The model’s prediction follows that of the DSV even though the effect of the DSV is unnoticeable.

### v2

@@ 4.2 Deep Support Vectors, From Dust to Diamonds
Firstly, the primal feasibility condition in Eq. (6) mandates support vectors be correctly classified by the trained model $\theta^{*}$. As presented in Sec. 4.1, instead of Eq. (3), we use a surrogate function for the loss:
$L_{\text{primal}}=\frac{1}{n}\sum_{i=1}^{n}L_{i},\quad\text{where}\quad L_{i}=\begin{cases}0&\text{if }\operatorname*{arg\,max}_{c}\Phi_{c}(x_{i};\theta^{*})=y_{i},\\ L(\Phi(x_{i};\theta^{*}),y_{i})&\text{otherwise}.\end{cases}$ (7)
This is designed to match the primal condition by mimicking the Hinge loss. When each DSV is correctly classified, no loss is incurred. Otherwise, we adjust the DSVs to align them with the correct target, effectively optimizing the support vectors. This approach also implicitly enforces the complementary slackness condition, as $L_{i}$ decreases confidence in the incorrect classification. Here, $L$ can be any loss function and we have employed the cross-entropy loss in our experiments.
@@
${\color[rgb]{0,0,0}L_{\text{stat}}=D(\theta^{*},-\sum_{i=1}^{n}\lambda_{i}\nabla_{\theta}L(\Phi(x_{i};\theta^{*}),y_{i}))}.$ (8)
@@
Moreover, as shown in Algorithm 1 (Appendix H), we satisfy the dual condition by ensuring the Lagrange multipliers $\lambda_{i}$’s are greater than zero and disqualify any $x_{i}$’s from being a support vector candidate if during optimization $\lambda_{i}$ becomes less than zero. The condition that is not explicitly satisfied is the complementary slackness. To directly fulfill the functional boundary for support vectors as specified in Sec. 3.2, we would need to be able to calculate the distance between functions, which is not only abstract but also requires a second-order computation cost. Therefore, we adopted a relaxed version of the KKT conditions that excludes this requirement. Furthermore, as demonstrated in Sec. 5.1, we have shown that DSVs implicitly satisfy the complementary slackness condition. This implies that DSVs meet every condition introduced in conventional SVM.
@@
As modern deep learning deals with extremely high-dimensional spaces, imposing additional constraints other than the primal feasibility and stationarity conditions is needed so that DSVs reside in the desired data manifold. To achieve this, we add a manifold condition, which enforces that the DSVs lie on the data manifold. By selecting DSVs that are in the intersection of the solution subspace and the data manifold, we can properly represent both the model and the training dataset.
To extract DSVs from the manifold, we assume that the model is well-trained, meaning it maintains consistent decisions despite data augmentation. In other words, the model should classify DSVs invariantly even after augmentation. To ensure this, we enforce that the augmented DSVs ($\mathcal{A}(x)$ where $\mathcal{A}$ denotes augmentation function) also meet the primary and stationarity conditions.
Also, we exploit traditional image prior [33, 16], total variance $L_{\text{tot}}$ and size of the norm $L_{norm}$ to make DSVs lie in the data manifold. $L_{\text{tot}}$ is calculated by summing the differences in brightness between neighboring pixels, reducing unnecessary noise in an image, and maintaining a natural appearance. $L_{norm}$, taking a similar role, penalizes the outlier and preserves important pixels.
Finally our DSV is obtained as follows where $\mathbb{E}_{\mathcal{A}}$ represents expectation over augmentations:
$\text{DSV}=\operatorname*{arg\,min}_{x}\mathbb{E}_{\mathcal{A}}\left[L_{\text{stationarity}}(\mathcal{A}(x))+\beta_{1}L_{\text{primal}}(\mathcal{A}(x))+\beta_{2}L_{\text{tot}}(x)+\beta_{3}L_{\text{norm}}(x)\right].$ (9)
One might wonder if there is a better sampling strategy than using DSVs, such as sampling far from the decision boundary instead of near it. We argue that in a high-dimensional data manifold, most data points are located close to the decision boundary because, in a data-scarce, high-dimensional space, every sample matters and thus would serve as a DSV.
@@ 5 Experiments
5.1 DSVs: Revival of Support Vectors in Deep Learning
DSVs meet SVM characteristics
@@ 5.1 DSVs: Revival of Support Vectors in Deep Learning
As discussed in Section 3.2, the principle of complementary slackness within the KKT conditions suggests that support vectors should be situated on the decision boundary, implying that support vectors typically exhibit high uncertainty from a probabilistic perspective, i.e., they possess high entropy. While DeepKKT does not explicitly incorporate the complementary slackness condition due to computational costs and ambiguity, Fig. 2(a) suggests that DSVs implicitly fulfill this condition; During the training process, we observe an increase in the entropy of DSV candidates, hinting that the generated DSVs are close to the decision boundary.
(a) Entropy change of DSV candidates over time
(b) Correlation between classwise mean test accuracy and the sum of $\lambda$’s
Figure 1: Characteristics of DSVs
img/cls ratio (%) Random selected DSVs
50 1 46.16 $\pm$ 1.93 48.91 $\pm$ 0.90
10 0.2 30.08 $\pm$ 1.96 33.69 $\pm$ 2.05
1 0.02 14.26 $\pm$ 0.99 16.83 $\pm$ 0.29
Table 1: In coreset selection benchmarks using the CIFAR-10 dataset, the DeepKKT condition is used as the selection criterion. Images with the highest $\lambda$ for each class were chosen to train a network.
In addition, we can infer the importance of a sample in the decision process by utilizing the DeepKKT condition. We trained the Lagrangian multiplier $\lambda$ for each test image. Figure 2(b) shows a strong correlation between the sum of $\lambda$ values for each class and its test accuracy. This finding is intriguing because, despite the model achieving nearly 100% accuracy during training due to overparameterization, DSVs provide insights into categorical generalization in the test phase. Not only does measuring its credibility indicate that a large $\lambda$ refers to an ‘important’ image for training, but $\lambda$ could also serve as a natural core-set selection measure. Table 1 shows this to be true. On the CIFAR-10 [11] dataset, we selected images with high $\lambda$ values and retrained the network with the selected images. In this case, the selected DSVs show higher test acccuracies compared to random selection. This characteristic resembles that of support vectors, as the model can be reconstructed with support vectors.
Finally, Fig. 1 demonstrates the high fidelity of the generated DSVs, providing practical evidence that these DSVs lie on the data manifold. Similar to how support vectors are reconstructed in an SVM, the DeepKKT condition enables the reconstruction of these vectors without referencing training data. This shows the effectiveness and adaptability of our DeepKKT in capturing key data features.
DSVs for Few Shot Dataset Distillation
DeepKKT emerges as a pioneering algorithm tailored for practical dataset distillation. DSVs addresses two critical concerns: 1) Protecting private information through data synthesis, and 2) Reducing the communication load by minimizing the size of data transmission. Traditional distillation algorithms encounter a fundamental paradox; as data predominantly originate from edge devices like smartphones [18, 17, 10], the requirement to access the entire dataset introduces significant communication overhead and heightens privacy concerns.
Figure 2: Model predictions for original versus DSV-informed edited images. (Top) Images were altered manually based on decision criteria derived from DSVs, influencing the model’s prediction. (Bottom) Images were altered based on DeepKKT loss.
Our DeepKKT relies solely on a pre-trained model without relying on the training dataset. This unique approach eliminates the need for edge devices to store or process large volumes of private data. As shown in Table 2, while traditional methods suffer significant performance drops under these scenarios and are incapable of implementing zero-shot scenarios. Conversely DeepKKT remains effective, requiring only minimal data: a single image per sample (i.e., initialization with real data), or in some cases, no images at all. For the zero-image setting, we initialized the images with data from other datasets to ensure diversity.
DSVs Encode the Decision Criteria Visually
Our findings suggest that DSVs not only satisfy the conditions of classical support vectors but also offers a global explanation of visual information. Fig. 2 experimentally verifies our claims and illustrates the practical use of DSVs, e.g., analysis of Fig. 1-Cifar10 reveals the decision criteria for classifying deer, cats, and dogs: 1) DSVs highlight antlers in deer, signifying them as a distinctive characteristic. 2) Pointed triangular ears are a recurring feature in DSVs of cats. 3) For dogs, a trio of facial dots holds significant importance. Using these observations, we altered a deer’s image by erasing its antlers and reshaping its ears to a pointed contour, which reduced the model’s confidence in classifying it as a deer, and caused the model to misclassify it as a cat. Similarly, by smoothing the ears of a cat image to diminish its classification confidence and then adding antlers or three facial dots, we influenced the model to reclassify the image as a deer or a dog, respectively. Additionally, the DeepKKT-altering case in Fig. 2-Bottom supports our assertions. Altering a flower image to resemble a deer class by changing the target class in the primal and dual feasibility loss, antlers grew similar to our manipulation.
This discovery holds significant implications about making models responsible; it introduces a qualitative aspect to assessing model performance. Consider a deer classification problem again. The model in our study would be less suitable, as evidenced by Fig. 1-Cifar10-deer, which indicates the model’s reliance on antlers for identifying deer – a feature not present in winter. This shows that DSVs enable us to conduct causal predictions by qualitatively analyzing models, as SVM does.
img/cls shot/class ratio (%) DC [37] DSA [35] DM [36] DSVs
1 0 0 - - - 21.68 $\pm$ 0.80
1 0.02 16.48$\pm$0.81 15.41$\pm$1.91 13.03$\pm$0.15 22.69 $\pm$ 0.38
10 0.2 19.66$\pm$0.78 21.15$\pm$0.58 22.42$\pm$0.43 -
50 1 25.90$\pm$0.62 26.01$\pm$0.70 24.42$\pm$0.29 -
500 10 28.06$\pm$0.61 28.20$\pm$0.63 25.06$\pm$1.20 -
10 0 0 - - - 30.35 $\pm$ 0.99
10 0.2 25.06$\pm$1.20 26.67$\pm$1.04 29.77$\pm$0.66 37.90 $\pm$ 1.69
50 1 36.44$\pm$0.52 36.63$\pm$0.52 36.63$\pm$0.52 -
500 10 43.55$\pm$0.50 44.66$\pm$0.59 47.96$\pm$0.95 -
50 0 0 - - - 39.35 $\pm$ 0.54
50 1 41.22$\pm$0.90 41.29$\pm$0.45 48.93$\pm$0.92 53.56 $\pm$ 0.73
500 10 52.00$\pm$0.59 52.19$\pm$0.53 60.59$\pm$0.41 -
Table 2: Performance of Few-shot learning on CIFAR10. ‘img/cls’ and ‘shot/class’ refer to the per-class number of generated images and the training samples used in generating the distilled dataset, respectively. ‘ratio’ is the ratio of the seen samples among the entire training samples. 0 shot refers to the distillation task performed without any access to the training data.
Figure 3: DSVs generated by a model that underwent transfer learning from CIFAR-10 to SVHN. During transfer learning, only the last layer is updated by SVHN.
5.2 Unlocking the potential of classifier as generator with DeepKKT
Practical Model Inversion with DeepKKT
In cloud environments or APIs, models are deployed with the belief that although they are sometimes trained with sensitive information, their black-box nature prevents users from inferring the data. This belief makes it possible to deploy sensitive models. However, as demonstrated in Fig 1, this belief is no longer valid. Fig. 3 further illustrates that model inversion remains feasible even in practical scenarios such as transfer learning scenarios, where only specific layers of a foundation model are fine-tuned. Remarkably, DeepKKT conditions enable model inversion even in these challenging environments, suggesting that they can be applied to a subset of the parameter space rather than the entire parameter space i.e., a more relaxed condition.
Classifier as Latent generative model
Considering the impressive capabilities of DSVs in the image generation domain, and the geometric interpretation that DSVs are samples near the decision boundaries, it is noted that enforcing the DeepKKT condition resembles the diffusion process. In each iteration, the DSV develops through the DeepKKT condition as follows:
$x_{t+1}=x_{t}-{\color[rgb]{0,0,0}\eta}\cdot\left([\nabla_{x}L_{\text{stat}}(\mathcal{A}(x_{t}))+\beta_{2}L_{\text{tot}}(x_{t})+\beta_{3}L_{\text{norm}}(x_{t})]+\beta_{1}\nabla_{x}L_{\text{primal}}(\mathcal{A}(x_{t}))\right).$ (10)
This is similar to the generalized form of the score-based diffusion process:
$x_{t+1}=x_{t}+\epsilon_{t}\cdot\left(\nabla_{x}\log p(x_{t})+\gamma\nabla_{x}\log p(y|x_{t})\right).$ (11)
Figure 4: Results showing DeepKKT images created solely by the primal condition or by the stationary condition. A sole usage of the primal condition shows low fidelity.
The first three loss terms in Eq. (10) aim to maximize the score ($\nabla_{x}\log p(x_{t})$), while the last term, the primal feasibility term, corresponds to the guidance term ($\gamma\nabla_{x}\log p(y|x_{t})$). As shown in Fig. 4, when only the primal loss term is used, meaningful DSV samples are not generated. This indicates that the other losses (stationarity and manifold terms) function update the image towards manifold i.e., score function. From this perspective, an arbitrarily assigned label $y$ can be used as a latent variable for guidance.
Figure 5: (Top) Generated DSVs using soft labels: $\delta$ was set 0.6, i.e., soft label $y=0.4y_{\text{left}}+0.6y_{\text{right}}$. (Middle) Examples of latent (soft-label) interpolation. (Bottom) Image Editing through latent.
To experimentally verify this, we performed a latent interpolation task and image editing, which is common in generative models [8, 3] By mixing different labels ($y_{i}=(1-\delta)y_{a}+\delta y_{b}$, where $y_{a}\neq y_{b}$), we generated DSVs as depicted in Fig. 5. When generating DSVs with these mixed soft labels, the generated DSVs semantically represent the midpoint between the two classes. A generated DSV either simply contains both images (the case of “balloon" and “spaniel") or semantically ‘fuse’ objects (the case of “lobster" and “harp", producing an image of a harp made out of lobster claws). For the image editing task, we assigned the latent variable to the desired class and then aligned the image using DeepKKT loss. The result was quite surprising: the method successfully transferred the image to the desired class while maintaining the original structure. For example, the sail of a yacht was seamlessly transformed into the shape of a scarf. This task was impossible with other methods; in diffusion models, for instance, a mask would be needed to edit the image seamlessly.
The fact that the generated images correctly merge the semantics of the classes suggests a couple of significant implications: 1) 1) New Generative Model: This approach offers a new type of generative model as an alternative to GANs and diffusion models. It can handle the same task without the need for training a specific model, as it leverages existing classification models for generative purposes. Furthermore, it is lightweight compared to diffusion models. For example, as the model size of a pretrained ResNet50 for ImageNet is only one-twentieth of that for SDXL [25], DSVs show a potential to leverage existing classification models for generative purposes. 2) Exploration of Classification Model Generalization: Unlike other generative models, classification models are trained simply to predict the label of an image. Yet, in latent interpolation and editing tasks, they demonstrate an understanding of semantics. This implies that, despite being trained to memorize class labels, the models grasp the overall semantics of the dataset. As they can generate seemingly unseen samples by interpolation and editing.
6 Conclusion
In this paper, we redefined support vectors in nonlinear deep learning models through the introduction of Deep Support Vectors (DSVs). We demonstrated the feasibility of generating DSVs using only a pretrained model, without accessing to the training dataset. To achieve this, we extended the KKT (Karush-Kuhn-Tucker) conditions to DeepKKT conditions and the proposed method can be applied to any deep learning models.
Akin to SVMs, the DeepKKT condition effectively encodes the decision boundary into DSVs. DSVs can reconstruct the model, making them useful for dataset distillation. Additionally, their visual encoding of the decision criteria can serve as a global explanation, helping to understand the model’s overall behavior and decisions. Furthermore, the DeepKKT condition transforms a classification model into a generative model with high fidelity. Not only can it sample data, but it also generalizes well, allowing the use of labels as latent variables.
Appendix A Societal Impacts
Our paper is closely related to Responsible AI (RAI), especially in enabling qualitative assessments of models. Our approach provides visual and intuitive explanations of a model’s decision-making criteria, offering insights that are both explanatory and responsible. Our approach of utilizing DSVs for RAI enables global explanations, surpassing traditional Explainable AI (XAI) methodologies, which usually focus on local explanations for individual inputs and cannot provide a global decision criterion. Furthermore, since our method is based on model inversion, it ensures safety and privacy. While the synthesized sets in Fig. 6 might appear similar to the selected sets, they do not replicate specific sample features. This is because DSVs represent a more generalized decision boundary, avoiding the inclusion of image-specific features. Consequently, DSVs enable all models using logistic loss to be more responsible.
Figure 6: Comparison of synthesized images (first row) created using the DeepKKT condition initiated from noise, and selected images (second row) from the CIFAR-10 training dataset. The selected images were chosen based on $\lambda$ values, i.e., each image has the highest $\lambda$ in each class. Both synthesized and selected images demonstrate similarity at the pixel level sharing common features.
Appendix B Limitations and Future work
In this paper, we propose the DeepKKT condition, which can be applied universally to any deep models to generate deep support vectors (DSVs) that function similarly to support vectors in SVMs. However, it should be noted that the equivalence between DSVs in deep learning models and support vectors in SVMs is only described intuitively, not rigorously. We have shown experimentally in Fig. 6 and intuitively in Sec. C why the DeepKKT condition should be as we suggested, but we have not derived it with rigorous math. Proving this rigorously would be a meaningful research topic.
Appendix C Intutive explanation of DeepKKT condition
Figure 7: The stationarity condition with a logistic loss. Even though the direction of the gradient $\hat{g}$ converges, the size of the gradient does not go to zero. Therefore, the direction of the converged gradient weight $\hat{w_{\infty}}$ aligns with $\hat{g}$.
In DeepKKT, many conditions make sense, except for one. For instance, the primal feasibility condition and the manifold condition are reasonable, and the dual feasibility condition can be regarded as importance sampling. However, the most counterintuitive part is the stationarity condition:
${\color[rgb]{0,0,0}L_{\text{stat}}=D(\theta^{*},-\sum_{i=1}^{n}\lambda_{i}\nabla_{\theta}L(\Phi(x_{i};\theta^{*}),y_{i}))}$ (12)
In this section, we will explain the dynamics of DSVs in an overparameterized deep network and how it is connected to deep learning. Below is a quick analogy of [28] to illustrate this connection.
A deep learning model follows the following ODE:
$w_{t+1}=w_{t}-\eta\nabla L(x,y;w_{t}).$ (13)
Here, $\eta$ is the learning rate and $t$ is the optimization step. The loss $L$ does not go to zero since deep learning models usually exploit a loss function with a logistic tail, such as the cross-entropy loss, and the gradient of the least confident sample (support vector) dominates overall gradient. Thus, there exists a convergence of the gradient direction $g_{\infty}\coloneqq\hat{\nabla}L$. There also exists a time $T$ where the gradient direction converges to $g_{\infty}-\varepsilon$ for a sufficiently small $\varepsilon$. As illustrated in Fig. 7, $w$ moves toward the direction of $-g_{\infty}$. Therefore, $\hat{w}_{\infty}\approx-g_{\infty}$.
This is for what stationarity condition wants to seek. The direction of $g_{\infty}$, by using only a few support vectors.
Appendix D Implementation Details
To obtain the results in Table 2 and Fig. 3, the ConvNet architecture [5] was used for pretraining $\Phi(\cdot;\theta)$ on the SVHN dataset [20], a digit dataset with dimensions similar to CIFAR-10 [11]. For ImageNet, we used the ResNet50 model [7] with the original setting in the paper. Specifically, we used the pretrained model in torchvision library in pytorch [22]. For visualizing synthesized DSVs in ImageNet, we increased the contrast in 224x224 dimensions. When calculating $L_{\text{stat}}$, we averaged the distance per parameter. In Alg. 1, $\eta$ was set to 5.
To synthesize DSVs in ImageNet, we used translation, crop, cutout, flip, and noise for augmentation, with hyperparameters set to 0.125, 0.2, 0.15, 0.5, and 0.01, respectively. In Eq. (9), we set $\alpha$ to 2e-5, $\beta$ to 40, and $\gamma$ to 1e-6. When calculating $L_{\text{stationarity}}$, we averaged the distance per parameter.
For dataset distillation in Table 2, we used translation, crop, flip, and noise for augmentation, with hyperparameters set to 0.125, 0.2, and 0.5, respectively. In Eq. (9), we set $\alpha$ to 2e-3, and both $\beta$ and $\gamma$ to 0. For retraining models with synthesized images, we used a learning rate of 1e-4 while the other parameters set to the default values of the Adam optimizer [9].
To obtain the pretrained weight $\theta^{*}$ for CIFAR10 and CIFAR100, we chose the ConvNet architecture [5], a common choice in deep learning. This architecture includes sequential convolutional layers followed by max pooling, and a single fully-connected layer for classification. The learning rate was set to $10^{-3}$ with a weight decay of $0.005$ using the Adam optimizer. Additionally, we employed flipping and cropping techniques, with settings differing from those used for DSVs reconstruction to ensure fair comparison. For pretraining $\Phi$ on the Street View House Numbers (SVHN) dataset [20], a digit dataset with dimensions similar to CIFAR-10 [11], we exclusively trained the fully-connected layer of the CIFAR-10 pre-trained ConvNet. This approach resulted in a training accuracy of 80%.
Appendix E DSVs by Selection
Fig. 8 shows the selected images with large Lagrangian multipliers $\lambda$’s, which correspond to the candidates used in Fig. 2(b). Surprisingly, there is a meaningful match between the selected DSVs and the synthesized DSVs in the CIFAR-10 dataset, as shown in Fig. 6. This implies that synthesizing DSVs corresponds to reviving training data that lie on the boundary manifolds.
Figure 8: Images of DSV candidates (Selected in the CIFAR-10 dataset).
Figure 9: Examples of adversarial attack in the CIFAR-10 dataset
Figure 10: More examples of latent interpolation in the ImageNet dataset
Figure 11: More examples of generated images with CIFAR-100 dataset
Figure 12: More examples of generated images with ImageNet dataset
Figure 13: More examples of image editing. The images to the left of the arrows represent the initial images before training, while those to the right depict the edited images after training.
Figure 14: More examples of image editing. The images to the left of the arrows represent the initial images before training, while those to the right depict the edited images after training.
Appendix F More Characteristics of DSVs